Two trains travel on tracks that run parallel to each other. Their starting points are 3740 miles apart. The first train, the Silver Streak, travels from northwest to southeast at 90 mph. The second train, the Red Blaze, travels from southeast to northwest at 80 mph. They start their journeys at the same time-midnight EST-and move toward each other. Assume that they travel at steady speeds without stops and without interruption. At what time will the trains meet? How far will the Silver Streak have traveled when they meet? How far will the Red Blaze have traveled when they meet?
Answer by rfer(16322)

moving together at 170 mph
3740/170=22 hrs
midnight plus 22 hrs= 10 pm
train SS traveled 22*90=1980
train RB traveled 22*80=1760